PLEAASEE HELP!! A battery with an emf of 12.0 V shows a terminal voltage of 11.6 V when operating in a circuit with two lightbulbs, each rated at 4.0 W (at 12.0 V),which are connected in parallel.

What is the battery's internal resistance?
Answer #1

Resistance, R = V^2/P = 12^2/4 = 36 ohm

Equivalent resistance of two bulbs = 36/2 = 18 ohm

Current, I = 11.6/18 = 0.64 A

Voltage drop, V' = 12 - 11.6 = 0.4 V

Internal resistance, r = 0.4/0.64 = 0.625 ohm

Comment in case any doubt please rate my answer...